### v5

# 1] organization=School of Computer Sciences , addressline=Universiti Sains Malaysia, city=11800 USM Penang, country=Malaysia

Abstract
Despite the remarkable results of deep learning in breast cancer histopathology image classification, challenges such as data imbalance and interpretability still exist and require cross-domain knowledge and collaboration among medical experts. This study proposes a breast cancer classification method using a dual-activated lightweight attention ResNet50 model, effectively addressing data imbalance and interpretability challenges. The model fuses a pre-trained deep ResNet50 and a lightweight attention mechanism to accomplish classification by embedding an attention module in layer 4 of ResNet50 and adding two fully connected layers. The fully connected network design employs LeakyReLU and ReLU activation functions.
The model outperforms SEResNet50, DensNet121, VGG16, VGG16Inception, ViT, Swin-Transformer, Dinov2_Vitb14, and ResNet50 models regarding precision, accuracy, recall, F1 score, and GMean, especially in the application performance on the BreakHis dataset. In particular, the model demonstrates significant robustness and broad applicability when dealing with the unbalanced breast cancer dataset. The model has been evaluated on histopathology images at magnification factors of 40X, 100X, 200X, and 400X, achieving accuracies of 98.5%, 98.7%, 97.9%, and 94.3%, respectively. The study comprehensively assessed the model’s performance. In the later stages of training, the validated losses and accuracies change minimally, showing that the model avoids overfitting and exhibits good generalization ability. This model exhibited the fastest convergence in all laboratory experiments, even though its parameters are not the smallest. This highlights the model’s efficacy as a lightweight attention framework, showcasing its efficiency in achieving rapid convergence without compromising performance.
keywords:
[…]
1 Introduction
Breast cancer, particularly among women, has emerged as one of the leading causes of fatal diseases, with its incidence on a continuous rise, eliciting worldwide health concerns [1, 2]. Early and accurate diagnosis and treatment are crucial in reducing the mortality rate of breast cancer. However, medical practices grapple with several obstacles, including inadequate healthcare resources and human error in interpretation, which hampers accurate and prompt detection and diagnosis of breast cancer. In this context, intelligent assisted diagnosis technology has become an innovative solution in breast cancer. Recent advancements in deep learning have had a transformative impact on breast cancer diagnosis. Deep learning theory has not only surpassed the performance limitations of traditional classification and recognition algorithms but has also integrated the entire feature extraction, feature selection, and classification process into a single step. Deep learning has simplified the training process by automatically extracting and classifying features through neural networks, eliminating redundancy and improving efficiency. These developments have revolutionized breast cancer diagnosis by improving accuracy and streamlining workflows [3]. This profound learning advancement has revolutionized breast cancer diagnosis [4, 5, 6, 7]. One key innovation in this area is developing and applying Convolutional Neural Networks (CNNs), a specific class of deep learning models designed to process data with a grid-like structure, such as medical images. The concept of CNN was first introduced back in 1998 [8]. The concept describes a method for efficiently classifying image data. CNNs gradually extract abstract features of an image through multi-level convolutional and pooling operations, allowing the network to extract necessary features from different categories of images so that the different categories can be classified accurately based on those features. CNNs have three main layers: convolutional, pooling, and fully connected. These layers work together to make CNNs a powerful tool for classifying medical images. The convolutional layer captures the image’s local features, the pooling layer reduces data dimensions while preserving significant information, and the fully connected layer maps these extracted features to various output classes, As shown in Fig. 1. However, despite the significant achievements of deep learning in medical image classification, conventional CNN architectures struggle with unbalanced medical image data. To tackle this issue, In this paper, a lightweight attention mechanism based on ResNet50 [9] is introduced to address the challenge of unbalanced medical image data in conventional CNN architectures. This novel approach effectively handles unstable category distributions and accurately detects malignant breast cancer in medical images. The critical contributions of the work can be summarized as follows:
(a) Proposing a lightweight attention mechanism network that adaptively adjusts channel feature extraction to improve feature representation.
(b) Integrating the attention mechanism module and ResNet50 to form a feature-rich model, the model achieves more efficient feature extraction based on fusing the two techniques. Specifically, in the fully connected network part of the attention mechanism module, we employed LeakyReLU [10] and ReLU [11] activation functions. Leaky-ReLU was chosen to mitigate the gradient vanishing problem, while ReLU was used to enhance the image feature representation, thus better capturing multi-scale spatial patterns.
(c) In this paper, test images with different magnifications and verify that the network is invariant to image scales and performs better than other state-of-the-art methods.
(d) The model was thoroughly evaluated by analyzing loss. The results show stable performance in the model’s later training phases, minimizing overfitting and exhibiting strong generalization for unfamiliar data.
This paper is organized as follows:
Section 2 provides an overview of the current state of the art in applying deep learning to breast cancer diagnostic techniques.
Section 3 details the proposed breast cancer classification method based on dual activation lightweight attention ResNet50.
Section 4 presents the dataset and implementation details of the experimental task and compares the classification results with other state-of-the-art networks.
Section 5 comprehensively evaluates the model’s performance through loss and accuracy analysis.
Finally, Section 6 provides a summary and conclusions.
Fig. 1: Overview of Convolutional Neural Network
2 Related works
Currently, two main methods are primarily used for cancer tissue histopathology image classification: traditional manual feature extraction and deep learning-based methods. Traditional methods rely on manually extracting image features and then using classifiers such as SVM or Random Forest for recognition; however, this method has shortcomings, such as a time-consuming feature extraction process, the requirement of domain knowledge, and the challenge of securing high-quality features [12]. In contrast, methods based on deep learning extract features through models like convolutional neural networks, which is more efficient [13]. Deep learning-based cancer tissue histopathology image classification primarily encompasses three methods: CNN-based breast cancer classification, transfer learning-based breast cancer classification, and attention mechanism-based breast cancer classification. The following section delves into these techniques in detail.
2.1 CNN
CNN [8], one of the most prevalent deep learning techniques, is apt for tasks involving image processing [14]. Through convolutional, pooling, and fully connected layers, CNNs are highly effective in learning image features and classifying them [15]. In breast cancer histopathology image classification, CNN models with various structures, such as AlexNet, Inception, VGGNet, DenseNet, ResNet, etc., can be trained on large datasets to learn features and classify breast cancer.
AlexNet, a deep convolutional neural network introduced by Alex et al. in 2012 [16], marked a pivotal breakthrough in the ImageNet image classification competition, setting a landmark in deep learning and CNNs. Omonigho et al. [17] employed a modified AlexNet with augmentation techniques on the MIAS mammography database for breast cancer detection, achieving an impressive 95.70% accuracy. This surpassed traditional methods, though parameter optimization and validation on unbalanced datasets require further exploration.
The Inception model is a neural network architecture introduced by Google, which aims to solve the computation problem and several parameters in neural networks and is suitable for image classification and related tasks. The hybrid model based on CNN-Inception-V4 was used by Nazir et al. [18]. It has excellent accuracy and sensitivity, but the model’s specificity is a little inferior, and it does not consider the processing computational efficiency and resource consumption issues.
VGGNet, a deep neural network architecture developed by researchers at Oxford University, performs well in image classification tasks. However, it is unsuitable for embedded devices or resource-constrained environments because of its high computational and memory requirements due to its considerable depth and number of parameters. Nedjar et al .[19] proposed a visualization method for explaining and understanding the decision-making process of VGG19 in classifying breast biopsy images, which helps to increase the trust and acceptance of medical professionals while providing a method for evaluating and Enhancing the efficiency of CNNs.
DenseNet is a deep neural network architecture developed by Gao Huang et al. [20]. It is known for its unique connectivity pattern. Each layer is densely connected to all previous layers, thus enabling the effectiveness of information flow and gradient propagation in deep networks. Jiménez Gaona et al. [21] introduced a novel approach to breast cancer detection and classification using the DenseNet architecture, which has advantages over other pre-trained CNN methods in solving problems such as gradient vanishing and reducing the number of parameters.
ResNet, a deep convolutional neural network architecture introduced by He et al. in 2015 [9], is particularly notable for its solution to the vanishing gradient issue in deep neural networks. This enables the development of intense neural networks, leading to substantial improvements in tasks like image categorization. Khikani et al. [22] proposed an enhanced capsule network using Res2Net blocks and additional convolutional layers to extract multi-scale features, and the results showed that their model outperformed previous deep-learning methods for breast cancer classification. However, ResNet-50 is prone to problems in handling unbalanced datasets. Given the significant superiority of the ResNet50 model over other methods in terms of accuracy, this paper has decided to modify the ResNet50 model to address these specific challenges.
2.2 Transfer learning
Transfer learning applies pre-trained models from one image domain to another, effectively bridging the gap between different image domains. In the context of breast cancer image classification, this pre-trained model approach involves fine-tuning CNN models, which were initially trained on a comprehensive large-scale image data set, and then fine-tuning it to adapt to the data set specifically used for breast cancer detection. Through this step, the model is optimized to better meet the specific needs of breast cancer classification. By using the inherent ability of the pre-trained model in generalization and feature extraction, this strategy significantly improves the efficiency of training and the accuracy of classification performance [23, 24, 25, 26, 27, 28, 29].
Ismail et al. [30] used transfer learning to expedite model training. Using models like VGG16 and ResNet50, they classified breast cancer images. Their experiments confirmed that transfer learning quickens model convergence and outperforms training from scratch.
Muhammad et al. [31] also utilized pre-trained deep learning models for breast cancer detection, such as ResNet50, ResNet101, VGG16, and VGG19. After training and testing these models on a dataset of 2453 breast histopathology images, they found that ResNet50 had a superior performance with an accuracy rate of 90.2%, an Area Under the Curve (AUC) of 90.0%, and a recall rate of 94.7%.
In conclusion, considering the exceptional robustness of ResNet50 in transfer learning, this study has chosen to further optimize and enhance it. However, we face specific challenges when dealing with imbalanced breast cancer datasets. Firstly, the imbalance in the dataset tends to bias the model towards categories with a larger number of instances, potentially leading to a significant reduction in the detection rate of key but less frequent categories, especially in early-stage breast cancer cases, thereby increasing the risk of missed diagnoses. Secondly, given the limited sample size of minority categories, there is a risk of overfitting in these samples, affecting the model’s generalization ability for novel and previously unseen minority class samples. Therefore, the goal of this study is to overcome these challenges by meticulously optimizing the application of ResNet50 in transfer learning, thereby enhancing the model’s overall performance in dealing with imbalanced breast cancer data.
2.3 Attention mechanism
In the field of medical imaging, the attention mechanism [32] in deep learning has proven to be a groundbreaking technique. It enables models to assign varying levels of importance to different regions within medical images, such as MRI scans, X-rays, CT scans, and histopathology images [33]. This targeted focus allows the models to prioritize and analyze the most pertinent areas for diagnosing diseases, identifying anomalies, or tracking the progression of medical conditions. By concentrating on crucial features within these images, the attention mechanism greatly enhances the accuracy and efficiency of automated medical image analysis, aiding in early detection and improved patient outcomes. In the context of breast cancer image classification, introducing the attention mechanism approach can prompt the model to learn critical and discriminative features in breast cancer images that are crucial for accurately distinguishing between malignant and benign tissues. This technique can improve the model’s performance while reducing the reliance on global image feature extraction and improving the accuracy of identifying critical breast cancer regions [34].
Researchers have proposed various breast cancer segmentation and classification systems incorporating attention mechanisms. For instance, Vardhan et al. [35] developed a breast cancer segmentation system that uses convolutional neural networks with interpretable artificial intelligence and the attention mechanism to improve breast cancer detection and classification accuracy and speed. Xu et al. [36] designed a system for breast cancer segmentation that uses a modified DenseNet architecture with an attention mechanism to optimize the convolutional neural network on the BreakHis dataset for high-precision classification of breast cancer. The method was tested on the BreakHis dataset, and the method’s performance was evaluated using accuracy metrics for both binary and multi-classification. Wang et al. [37] proposed a breast cancer classification model that combines deep migration learning and visual attention mechanisms to improve the accuracy and efficiency of real-world clinical diagnosis. This deep-migration-based attention mechanisms approach uses migration learning to overcome the overfitting problem and combines the attention module to focus on characterizing breast lesions. Wang et al.’s work [37] achieves high-precision results that outperform other state-of-the-art breast image classification algorithms. This section summarizes recent scholarly research and exploration in breast cancer classification, focusing on the use of the attention mechanism. The insights gleaned indicate that applying the attention mechanism in breast cancer image classification methods can enhance the accuracy of crucial region identification and boost the interpretability of the model simultaneously. Still, the difficulty of training and parameter tuning is also increased accordingly. In this study, a lightweight design of the attention module is proposed, which incorporates adaptive average pooling, a two-layer fully connected network, and the dual activation function of LeakyReLU and ReLU to efficiently extract and enhance the critical information in the image through item-by-item multiplication (element-wise multiplication of individual pixels and features within the images)of features images, aiming to improve performance and reduce computational consumption.
3 Methods
3.1 Transfer ResNet50
The ResNet50 [9] model has a relatively complex structure, including a 50-layer deep network structure. The specific structure is shown in Table 1. The primary module is made up of numerous residual blocks. Each block is equipped with multiple convolutional layers and distinctive identity maps. Uniquely, the ResNet50 model introduces the concept of residual learning. This innovation solves the problem of vanishing gradients in deep networks, thereby opening up the possibility of building deeper neural networks than ever before. This design breakthrough allowed for achieving exceptional performance. However, considering the efficient utilization of data and computing resources, the transfer learning strategy is adopted in this paper based on the pre-trained ResNet50 model for model building. Through transfer learning, full benefit can be obtained from the high-level feature representations learned by ResNet50 on large-scale datasets in medical image classification tasks. This allows for swiftly attaining superior classification outcomes, even with constrained data and computing capacities.
Table 1: The ResNet50 Model
Module Name Output Size 50 Layers
Conv1 $112\times 112112 × 112$ $7\times 77 × 7$, 64 stride=2
[…]
Pooling $1\times 11 × 1$ Average Pool, 1000-d FC.
Note: The term "1000-d FC" in the pooling layer refers to a fully-connected layer consisting of 1000 neurons.
3.2 Dual-activated lightweight attention ResNet50
The paper proposes a dual-activation lightweight attention ResNet50 model method. Initially, a pre-trained ResNet50 is utilized as the foundational architecture. Building upon this, an innovative dual-activation lightweight attention module is introduced, along with a newly designed fully connected layer tailored for specific tasks. This attention module is integrated into the fourth residual block of ResNet50, characterized by the amalgamation of adaptive average pooling and a two-layer fully connected network, employing LeakyReLU[10] and ReLU [11] as its dual activation functions. Additionally, the newly added fully connected layer is positioned at the end of the modified ResNet50 model, primarily aimed at addressing the complexities of binary classification problems.
In the dual-activation lightweight attention module proposed in this paper, adopt the LeakyReLU activation function [10] for the initial fully connected layer, leveraging its competence in allowing for a slight gradient flow under zero. This is done to mitigate the issue of vanishing gradients, thereby enhancing the training efficacy of deep neural networks. The ReLU activation function [11] is chosen for the second fully connected network to further improve the model performance and image feature representation.
Combining these two activation functions greatly enhances the model’s expressive power and capacity to learn intricate image features, making the model more suitable for image classification tasks. In the experiments, this design improves accuracy and performance while maintaining the convergence speed of the model. The flowchart and visualization of the "Dual Activation Lightweight Attention ResNet50 Module" can be comprehended through the displayed in Fig. 2 and Fig. 3, respectively, and the overall technical framework is shown in Fig. 4.
Fig. 2: Flowchart for dual-activated lightweight attention ResNet50 model
Fig. 3: Visualization of dual-activated lightweight attention ResNet50 model
Fig. 4: The framework of dual-activated lightweight attention ResNet50 model
4 Experiment
4.1 BreakHis Dataset
The BreakHis dataset, utilized in this study focusing on breast cancer histopathological images, was generously provided by scholar Spanha [38]. This dataset comprises histopathology samples of breast cancer tumors obtained from 82 patients. The samples were sectioned under a microscope and then imaged at different magnifications (40×, 100×, 200×, 400×) to capture intricate details. Experienced medical professionals meticulously interpreted and identified the images. The dataset consists of a total of 7,909 breast cancer pathological images, comprising 2,429 benign tumor images and 5,429 malignant tumor images. Each image has dimensions of 700 × 460 pixels and is stored in PNG format with three color channels. The detailed statistics of the dataset can be found in Table 2, covering sample size, class distribution, etc. This multi-angle data resource provides a solid foundation for effective image classification and analysis research. In addition, to more intuitively display the breast cancer tissue details under different magnifications, Fig. 5 presents these pathological images as a practical example.
Table 2: BreakHis dataset: benign and malignant image distribution across magnification factors
Class Magnification Factor
40X 100X 200X 400X
Benign
625 644 623 588
Malignant
1370 1437 1390 1232
Total No of Images
1995 2081 2013 1820
Refer to caption 40x
Refer to caption 100x
Refer to caption 200x
Refer to caption 400x
Fig. 5: Images of breast cancer pathology at four magniations
4.2 Experimental environment
The experiments were carried out on an NVIDIA RTX A4000 GPU, utilizing the PyTorch 2.0.1 framework paired with Python 3.10.
4.3 Experiment implementation
This research utilizes a ratio of 6:1:3 to divide the processed dataset into training, verification, and test sets, as presented in Table 3. The training set is utilized for model training, during which parameter learning occurs. The verification set is employed to assess the model’s generalization ability. Through continuous verification and parameter fine-tuning, superior model performance is ensured, with the optimal model being retained during the verification process. The test set measures the model’s recognition rate and generalization performance. Before model training, the training data is shuffled to enhance the model’s robustness and generalization capabilities.
In deep learning, the configuration of various parameters is paramount. Particularly, the learning rate plays a crucial role; an excessively high learning rate may result in the loss value exploding, while a meager learning rate may lead to overfitting and reduce the model’s convergence speed. In previous studies, the initial learning rate is usually set in the range of 0.1 to 0.0001. Considering the relatively small size of the BreakHis medical imaging dataset (containing 7909 images) and the uneven distribution of the data, a lower learning rate of 0.0001 was chosen for this study so that the model could adjust the weights more accurately. This is important for dealing with small and unbalanced datasets to ensure that all categories are given proper attention. In addition, the setting of the training batch size and the number of iterations is extremely important. In addition, the setting of the training batch size and the number of iterations is extremely important. Previous studies usually set the batch size between 32 and 128. Considering the GPU’s 16GB video memory limit, the lowest batch size of 32 was chosen for this study. As for the number of iterations, it is also usually set between 50 and 128. Given the small size of the BreakHis dataset, this study limited the number of iterations to 50, aiming to achieve a balance between training effectiveness and computational efficiency. Finally, this study also specifically considered the setting of the dropout rate. Previous studies generally set the dropout rate between 0.25 and 0.5. In the processing of the unbalanced BreakHis dataset, a lower dropout rate helps to ensure that the categories with smaller numbers are sufficiently learned, thus improving the model’s recognition accuracy and generalization ability for these categories, thus this study chose a lower dropout rate of 0.25.
The method proposed in this study as well as the benchmark model illustrated in Table 5 apply the above settings to ensure the fairness of the experiments and the comparability of the results. These carefully selected parameters provide a solid foundation for the experimental design and ensure the reliability and efficiency of the model throughout the study, the parameter settings are shown in Table 4.
Table 3: Dataset split for experimental task
Split Percentage Purpose
Training 60% Model Training
Validation 10% Hyperparameter Tuning
Testing 30% Model Evaluation
Table 4: Experimental configuration
Parameter Value
Batch Size 32
Epochs 50
Learning Rate 0.0001
Dropout 0.25
4.4 Baseline selection
Table 5 presents a comparative analysis of various deep learning methods, encompassing both widely adopted approaches and the novel strategy introduced in this paper. The following section briefly outlines several of the models included in the comparison:
SEResNet50 [39]: Hu et al. proposed this model in 2018. The key feature of this model is integrating the "squeeze and excitation" attention mechanism, which aims to improve the network’s capabilities.
DenseNet-121 [40]: Huang et al. proposed this model in 2017 by introducing dense connections (dense blocks) to transfer and reuse information between network layers, thus improving the efficiency of feature transmission and reproduction. This study employs transfer learning with DenseNet-121 is employed for breast cancer classification. The advantage of using a pre-trained DenseNet is that it can adapt to small sample data.
VGG16 [41]: This model is from Simonyan and Zisserman and was released in 2014. It is known for its clean structure and thoughtful design philosophy and is widely used for image classification and feature extraction tasks, relying mainly on deeply stacked convolutional layers.
VGG16Inception [42]: The model is presented by M. Saini and S. Susan. The article delves into a deep learning model called "VGGIN-Net," mainly applied to analyze unbalanced breast cancer datasets. The study introduces a new deep learning framework that incorporates transfer learning to address the problem of category imbalance in biomedical datasets, applied explicitly to the BreakHis breast cancer dataset. To reduce training time and computational cost, the model uses a pre-trained network to migrate learned weights from one domain to another."
ViT (Vision Transformer) [43]: proposed by Dosovitskiy et al. in 2020, it has gained widespread attention by successfully applying the Transformer architecture to image recognition tasks. This model has shown great strength in image recognition.
Swin-Transformer [44]: Liu et al. created this model in 2021, and its performance in various vision tasks is outstanding. This model is unique in that it solves the computational and memory problems that can be encountered when processing large images.
Dinov2_Vitb14 [45]: proposed in 2023 by a team of researchers at Facebook, these models contain a pre-training dataset of 142 million unlabelled images to obtain efficient visual properties demonstrating general solid performance.
Each of the above models has uniqueness and excellent image classification and vision tasks. To emphasize the effectiveness of the proposed approach, an in-depth comparison of all these models has been conducted, achieving significant improvements in various evaluation metrics.
Table 5: Classification results of 9 network models
Network Models 40X 100X
Accuracy F1 IBA GMean Accuracy F1 IBA GMean
SEResNet50 [39]
0.883 0.779/0.883 0.860 0.860 0.850 0.720/0.849 0.813 0.812
DensNet121 [40]
0.801 0.639/0.800 0.763 0.762 0.811 0.654/0.809 0.755 0.753
VGG16 [41]
0.902 0.813/0.902 0.877 0.877 0.850 0.724/0.851 0.830 0.829
VGG16Inception [42]
0.977 0.955/0.977 0.965 0.965 0.971 0.943/0.971 0.964 0.964
ViT [43]
0.835 0.693/0.833 0.769 0.767 0.834 0.694 /0.833 0.788 0.787
Swin-Transformer [44]
0.843 0.711/0.844 0.819 0.818 0.877 0.767/0.876 0.836 0.835
Dinov2_Vitb14 [45]
0.873 0.762/0.873 0.843 0.843 0.869 0.753/0.868 0.835 0.834
RseNet50 [9]
0.927 0.860/0.927 0.914 0.914 0.944 0.891/0.944 0.922 0.922
Dual-Activated Lightweight Attention ResNet50 Model 0.985 0.970/0.985 0.980 0.980 0.987 0.974/0.987 0.985 0.985
Network Models 200X 400X
Accuracy F1 IBA GMean Accuracy F1 IBA GMean
SEResNet50 [39]
0.906 0.820/0.906 0.876 0.875 0.876 0.770/0.878 0.864 0.863
DensNet121 [40]
0.812 0.659/0.812 0.775 0.774 0.779 0.603/0.777 0.737 0.736
VGG16 [41]
0.851 0.722/0.850 0.812 0.811 0.848 0.718/0.848 0.817 0.817
VGG16Inception [42]
0.954 0.910/0.954 0.938 0.938 0.889 0.788/0.888 0.857 0.856
ViT [43]
0.841 0.705/0.840 0.801 0.800 0.832 0.691/0.831 0.803 0.803
Swin-Transformer [44]
0.876 0.766/0.875 0.838 0.837 0.846 0.724/0.851 0.791 0.789
Dinov2_Vitb14 [45]
0.853 0.726/0.852 0.802 0.800 0.885 0.782/0.885 0.840 0.839
RseNet50 [9]
0.960 0.922/0.960 0.947 0.947 0.925 0.856/0.925 0.918 0.918
Dual-Activated Lightweight Attention ResNet50 Model 0.979 0.958/0.979 0.98 0.98 0.943 0.891/0.944 0.945 0.945
Note: In the table, the DenseNet-121 model indicates that employed a pre-trained DenseNet-121 model for transfer learning.
Table 6: The count of parameters in the network model and the time it takes to converge.
Network Models Convergence Time (seconds) Total Parameters Average Time
40X 100X 200X 400X
SEResNet50 [39]
2696.22 2814.30 2711.89 2356.18 27073090 2644.65
DensNet121 [40]
2352.46 2643.79 2423.36 2078.34 9028970 2347.49
VGG16 [41]
3212.95 3182.54 3011.53 2640.19 134268738 3011.80
VGG16Inception [42]
3074.99 3253.50 3303.12 2682.61 38139938 3078.56
ViT [43]
2932.11 3022.54 3013.86 2691.16 51375106 2914.92
Swin-Transformer [44]
3020.11 3132.42 2940.98 2592.96 86745274 2921.62
Dinov2_Vitb14 [45]
2895.02 2303.25 2245.98 2017.15 86582018 2365.35
RseNet50 [9]
2920.53 2933.26 2307.94 2097.33 21797672 2564.77
Dual-Activated Lightweight Attention ResNet50 Model 2133.02 2319.50 1990.58 1896.82 25611842 2084.98
Note: In the table, the DenseNet-121 model indicates that employed a pre-trained DenseNet-121 model for transfer learning.
4.5 Experiment metrics
Used four metrics, Accuracy [46, 47, 48], F1 Score [49], IBA [50], and GMean [51], to assess the performance of the model holistically. The F1 Score is a crucial metric that combines accuracy and recall, providing a comprehensive measure of a model’s ability to classify positive and negative instances. This is especially useful for imbalanced datasets where positive and negative samples differ significantly. [52]. IBA (Imbalance Balanced Accuracy) takes a targeted approach to tackle the class imbalance challenge by harmonizing accuracy and balance. This approach offers a more holistic assessment of the model’s performance, encompassing sample diversity and fairness [53]. GMean (Geometric Mean) is a metric that calculates the geometric mean of individual classification accuracies, which is used to evaluate the model’s ability to handle category imbalance and ensure that the model can perform well across all classes [54].
In Table 5 showcases the performance of the model on the test set across various image resolutions, such as 40X, 100X, 200X, and 400X. Regardless of the solution, the model consistently outperforms nine other popular network models regarding accuracy, F1 score, IBA, and GMean. These results demonstrate that the network model has a significant advantage regardless of which image resolution is used. This gives us a deep insight into the model’s excellent performance and great potential in performing medical image classification tasks.
In Table 6, presents a comparison of the model with eight other network models regarding the number of model parameters and convergence speed. Notably, the DenseNet-121 model possesses the fewest parameters. However, the model has the shortest convergence time and the fastest average convergence speed across all image resolutions, including 40X, 100X, 200X, and 400X. This further reinforces the superior performance and potential of the model for medical image classification tasks. These results proved the efficiency and effectiveness of the model.
5 Performance evaluation and visualization
To thoroughly oversee the model’s performance and glean insights into the training process, focus on several key metrics, including training loss, validation loss, training accuracy, and validation accuracy. Visualizing these metrics can demonstrate the model’s training performance. The results of analyzing these key metrics are presented in detail below.
Refer to caption 40X
Refer to caption 100X
Refer to caption 200X
Refer to caption 400X
Fig. 6: Training and validation loss with error bars
5.1 Loss analysis
This study utilized a dual-activated lightweight attention ResNet50 model for breast cancer classification. The critical metrics for monitoring model performance and understanding the training process are training loss and validation loss [55, 56, 57]. These loss values play an important role when evaluating model performance, and their visualization gives us a clearer picture of the training progress.
Fig. 6 shows the variation of training loss and validation loss with the number of training iterations for different image resolutions (including 40X, 100X, 200X, and 400X). The x-axis denotes the quantity of training iterations, while the y-axis signifies the values of the loss function. Used line plots to show the trend of training loss and validation loss and added error bars to each data point to show the standard deviation of the loss values.
As training progresses, a gradual decrease in training loss and validation loss is observed, which signifies that the model is effectively learning and adapting to the data. This is a positive trend and shows that the model is making progress on the task. In the later stages of training, the validation loss changes less. It offers a stable trend, indicating that the model steers from overfitting on the validation set and demonstrates strong generalization to new data.
Refer to caption 40X
Refer to caption 100X
Refer to caption 200X
Refer to caption 400X
Fig. 7: Training and validation accuracy with error bars
5.2 Accuracy analysis
In Fig. 7, the evolution of training accuracy and validation accuracy at different image resolutions (40X, 100X, 200X, and 400X) is depicted [58]. The horizontal axis represents the number of training iterations, while the vertical axis represents the accuracy. A line graph is employed to portray the accuracy trend, and error bars are incorporated to convey the standard deviation of the accuracy values.
After meticulous analysis, the following outcomes have been derived: Training and validation accuracy undergo a gradual ascent, showcasing the model’s persistent learning and adaptation throughout the training process. This upward trajectory suggests that the model is advancing in comprehending and classifying the data. Additionally, the validation accuracy maintains stability, especially in the later stages of training, implying that the model avoids overfitting to the training data and can generalize effectively to unseen data. These findings are pivotal in comprehending the model’s performance and can facilitate further optimization for enhanced breast cancer classification results.
6 Summary and conclusions
This paper focuses on applying a lightweight attention mechanism to a ResNet50 model for classifying medical histopathology tissue images, mainly when dealing with limited-scale datasets. The ResNet50 model incorporates the lightweight attention mechanism and a dropout layer to optimize the pre-trained deep network layer. The performance evaluation includes metrics such as accuracy, F1 score, IBA, and GMean, which allow for a comprehensive assessment of this paper model’s performance compared to existing techniques. The experimental results highlight significant advantages, particularly when addressing the challenges of unbalanced breast cancer datasets. The model demonstrates improved robustness, generalization, faster convergence during training, and effective prevention of overfitting. These findings provide a solid basis for enhancing classification accuracy. In future research, investigating more breast multiclassification problems is planned.
[…]