### v10

Abstract
Automatic breast cancer classification in histopathology images is crucial for precise diagnosis and treatment planning. Recently, classification approaches based on the ResNet architecture have gained popularity for significantly improving accuracy by using skip connections to mitigate vanishing gradient problems, thereby integrating low-level and high-level feature information. Nevertheless, the conventional ResNet architecture faces challenges such as data imbalance and limited interpretability, necessitating cross-domain knowledge and collaboration among medical experts. This study effectively addresses these challenges by introducing a novel method for breast cancer classification, the Dual-Activated Lightweight Attention ResNet50 (DALAResNet50) model. It integrates a pre-trained ResNet50 model with a lightweight attention mechanism, embedding an attention module in the fourth layer of ResNet50 and incorporating two fully connected layers with LeakyReLU and ReLU activation functions to enhance feature learning capabilities. The DALAResNet50 method was tested on breast cancer histopathology images from the BreakHis Database across magnification factors of 40X, 100X, 200X, and 400X, achieving accuracies of 98.5%, 98.7%, 97.9%, and 94.3%, respectively. It was also tested on the four-class classification BACH dataset and the three-class classification Mini-DDSM dataset, where it also performed excellently. Additionally, the model was compared with established deep learning models such as SEResNet50, DenseNet121, VGG16, VGG16Inception, ViT, Swin-Transformer, Dinov2_Vitb14, and ResNet50. The reported results of DALAResNet50 have been shown to outperform the compared approaches regarding accuracy, F1 score, IBA, and GMean, demonstrating significant robustness and broad applicability when dealing with different magnifications and imbalanced breast cancer datasets, proving the model’s wide applicability. To address the interpretability challenges, we introduce the Dynamic Threshold Grad-CAM (DT Grad-CAM) method. DT Grad-CAM enhances the traditional Grad-CAM approach by applying adaptive thresholding using Otsu’s method, providing clearer and more focused visualizations of significant regions. This method improves the clarity and relevance of the highlighted areas, making it easier for medical experts to interpret the model’s predictions and understand the critical features contributing to the classification. The integration of DT Grad-CAM with DALAResNet50 not only ensures high classification performance but also offers enhanced interpretability, crucial for reliable and transparent breast cancer diagnosis.
Index Terms:
Lightweight attention mechanism, Convolutional Neural Networks, breast cancer classification.
[…]
=-21pt
I Introduction
Breast cancer has become a leading cause of mortality, particularly among women, with its incidence rising continuously, raising global health concerns [1, 2]. Early, accurate diagnosis and treatment are pivotal in decreasing the mortality rate associated with breast cancer. Yet, medical practices face challenges such as insufficient healthcare resources and human error in interpretation, which impede the timely and precise detection and diagnosis of breast cancer. Against this backdrop, intelligent diagnostic assistance technologies have emerged as innovative solutions in the battle against breast cancer. Recent advances in deep learning have revolutionized breast cancer diagnosis, surpassing the limitations of traditional classification and recognition algorithms by integrating feature extraction, selection, and classification into a unified step. Deep learning automates feature extraction and classification through neural networks, reducing redundancy and enhancing efficiency, thus significantly improving diagnostic accuracy and workflow [3, 4, 5, 6, 7]. A pivotal innovation in this domain is developing and applying Convolutional Neural Networks (CNNs), specialized deep learning models adept at processing data, such as medical images, with a grid-like topology. In 1998, LeCun et al. [8] introduced Convolutional Neural Networks (CNNs), which progressively extract abstract features from images through multiple levels of convolutional and pooling operations. This process enables CNNs to achieve classification based on these extracted features. CNNs comprise convolutional, pooling, and fully connected layers, collaboratively facilitating the robust classification of medical images. The convolutional layer identifies local image features; the pooling layer reduces data dimensions while retaining critical information; and the fully connected layer maps these features to output classes. Despite deep learning’s significant achievements in medical image classification, conventional CNN architectures often struggle with imbalanced medical image data. This paper proposed a Dual-Activated Lightweight Attention ResNet50 (DALAResNet50) model designed to overcome the challenge of imbalanced data in CNN architectures. This innovative approach efficiently manages uneven category distributions and enhances the detection of malignant breast cancer in medical images. The primary contributions of this work are summarized as follows:
(a) This study introduces a novel lightweight attention mechanism network that adaptively adjusts channel feature extraction to enhance feature representation. This method optimizes the network’s ability to recognize features relevant to breast cancer classification. Performance was evaluated on the BreakHis dataset for binary classification, the BACH dataset for four-class classification, and the Mini-DDSM dataset for three-class classification. Evaluation metrics included accuracy, F1 score, IBA, and GMean, which allow for a comprehensive assessment of the model’s performance compared to existing techniques. Experimental results highlighted significant advantages, especially in addressing the challenges of imbalanced breast cancer datasets. The model demonstrated improved robustness, generalization, faster convergence during training, and effective prevention of overfitting.
(b) This paper introduces a DT Grad-CAM method to address the interpretability challenges commonly associated with deep learning models. DT Grad-CAM enhances the traditional Grad-CAM by applying adaptive thresholding using Otsu’s method, providing clearer and more focused visualizations of important regions. This approach improves the clarity and relative prominence of highlighted areas, making it easier for medical experts to interpret the model’s predictions and understand the key features contributing to the classification.
Organization of the paper:
Section II reviews the latest advancements in applying deep learning for breast cancer diagnosis.
Section III proposed the DALAResNet50 model approach for classifying breast cancer histopathology images.
Section IV outlines the dataset and implementation specifics of the experimental setup and compares the classification outcomes with those of other leading networks.
Section V conducts an in-depth performance evaluation of the model through loss and accuracy analysis.
Section VI provides a summary and conclusions.
II Related works
Cancer tissue histopathology image classification predominantly utilizes traditional manual feature extraction and deep learning-based methods. Traditional approaches depend on manually extracting features from images and employing classifiers such as SVM (Support Vector Machine) or Random Forest for recognition. However, these methods encounter limitations, such as the need for intensive feature extraction, a requirement for substantial domain knowledge, and challenges in acquiring high-quality features [12]. In contrast, deep learning-based methods, mainly through models like Convolutional Neural Networks (CNNs), offer a more efficient feature extraction process [13]. Specifically, deep learning approaches to cancer tissue histopathology image classification include CNN-based, transfer learning-based, and attention mechanism-based breast cancer classification. Each of these methods leverages the inherent strengths of deep learning to improve classification performance, which will be explored in detail in the subsequent section.
II-A CNN
CNN, one of the most widespread deep learning techniques, excels in image processing tasks [14]. They utilize convolutional, pooling, and fully connected layers to learn and classify image features effectively [15]. In breast cancer histopathology image classification, CNN models like AlexNet, Inception, VGGNet, DenseNet, and ResNet have demonstrated substantial capabilities when trained on large datasets.
AlexNet, introduced in 2012 by Alex et al. [16], represented a significant milestone in ImageNet competition history, showcasing the potent capabilities of CNNs. Modified versions of AlexNet, such as the one by Omonigho et al. [17] on the MIAS mammography database, achieved notable accuracy, indicating superior performance over traditional methods. However, challenges remain in parameter optimization and application to imbalanced datasets.
The Inception model, developed by Google, addresses computational and parameter efficiency in neural networks, making it ideal for image classification tasks. Nazir et al.’s hybrid CNN-Inception-V4 model [18] demonstrated high accuracy and sensitivity, though it faced limitations in specificity and computational efficiency considerations.
VGGNet, crafted by Oxford University researchers, excels in image classification but is less suited for resource-constrained environments due to its intensive computational and memory demands. Nedjar et al. [19] presented a visualization method for VGG19’s decision-making process in breast biopsy image classification, enhancing medical professionals’ trust and CNN efficiency.
DenseNet [20], known for its efficient information and gradient flow through dense connectivity, offers advantages in overcoming challenges like gradient vanishing. Jiménez Gaona et al. [21] leveraged DenseNet for breast cancer detection, showing its superiority in reducing parameters and improving performance compared to other CNN architectures.
ResNet, introduced by He et al. in 2015 [9], addresses deep neural networks’ vanishing gradient problem, allowing for the construction of deeper networks. An enhanced capsule network using Res2Net blocks, proposed by Khikani et al. [22], demonstrated improved performance in breast cancer classification, although ResNet-50 faces challenges with imbalanced datasets. This paper presents modifications to the ResNet50 model to tackle these issues, given its accuracy advantage over other methods.
II-B Transfer learning
Transfer learning applies pre-trained models from one image domain to another, bridging the gap between different image domains. In the context of breast cancer image classification, this pre-trained model approach involves fine-tuning CNN models, which were initially trained on a comprehensive large-scale image dataset, and then fine-tuning it to adapt to the dataset used explicitly for breast cancer detection. This step optimizes the model better to meet the specific needs of breast cancer classification. By using the inherent ability of the pre-trained model in generalization and feature extraction, this strategy significantly improves the efficiency of training and the accuracy of classification performance [23, 24, 25, 26, 27, 28, 29].
Ismail et al. [30] used transfer learning to expedite model training. Using models like VGG16 and ResNet50, they classified breast cancer images. Their experiments confirmed that transfer learning quickens model convergence and outperforms training from scratch.
Muhammad et al. [31] also utilized pre-trained deep learning models for breast cancer detection, such as ResNet50, ResNet101, VGG16, and VGG19. After training and testing these models on a dataset of 2453 breast histopathology images, they found that ResNet50 had a superior performance with an accuracy rate of 90.2%, an Area Under the Curve (AUC) of 90.0%, and a recall rate of 94.7%.
In conclusion, this study has focused on further optimization and enhancement due to the exceptional robustness of ResNet50 in transfer learning. Nevertheless, the investigation encounters specific challenges when dealing with imbalanced breast cancer datasets. Firstly, the imbalance in the dataset tends to bias the model towards categories with a larger number of instances, potentially leading to a significant reduction in the detection rate of key but less frequent categories, especially in early-stage breast cancer cases, thereby increasing the risk of missed diagnoses. Secondly, given the limited sample size of minority categories, there is a risk of overfitting in these samples, affecting the model’s generalization ability for novel and previously unseen minority class samples. Therefore, this study aims to overcome these challenges by meticulously optimizing the application of ResNet50 in transfer learning, thereby enhancing the model’s overall performance in dealing with imbalanced breast cancer data.
II-C Attention mechanism
In medical imaging, the attention mechanism [32] in deep learning has emerged as a revolutionary technique. It allows models to allocate varying levels of importance to different regions within medical images, such as MRI scans, X-rays, CT scans, and histopathology images [33]. This selective focus enables the models to prioritize and scrutinize the most relevant areas for diagnosing diseases, detecting anomalies, or monitoring the progression of medical conditions. By zeroing in on essential features within these images, the attention mechanism significantly boosts the accuracy and efficiency of automated medical image analysis, contributing to early detection and enhanced patient outcomes. Within breast cancer image classification, incorporating the attention mechanism encourages the model to discern critical and distinctive features in breast cancer images that are pivotal for accurately differentiating between malignant and benign tissues. This strategy can elevate the model’s performance, diminish reliance on global image feature extraction, and enhance the precision in pinpointing vital breast cancer regions [34].
Researchers have proposed various breast cancer segmentation and classification systems incorporating attention mechanisms. For example, Vardhan et al. [35] developed a breast cancer segmentation system utilizing Convolutional Neural Networks (CNNs) with interpretable artificial intelligence and the attention mechanism to enhance breast cancer detection and classification accuracy and speed. Xu et al. [36] designed a system for breast cancer segmentation employing a modified DenseNet architecture with an attention mechanism to refine the CNN’s performance on the BreakHis dataset for high-precision breast cancer classification. This method was evaluated on the BreakHis dataset using accuracy metrics for both binary and multi-classification. Wang et al. [37] proposed a breast cancer classification model that combines deep transfer learning and visual attention mechanisms to increase the accuracy and efficiency of real-world clinical diagnoses. Based on deep transfer learning with attention mechanisms, this approach utilizes transfer learning to tackle the overfitting problem and integrates an attention module to concentrate on characterizing breast lesions. Wang et al.’s work achieves high-precision results that surpass other state-of-the-art breast image classification algorithms.
This section summarizes recent scholarly research and exploration of breast cancer classification, emphasizing the use of the attention mechanism. The insights indicate that applying the attention mechanism in breast cancer images and classification methods can simultaneously enhance the accuracy of crucial region identification and increase the model’s interpretability. However, this also correspondingly increases the following difficulties:
II-C1 Limitations in global feature extraction
Although the attention mechanism is designed to improve model performance by focusing on critical sections of the image, thereby playing a positive role in enhancing the model’s ability to process global information of the image, it may still encounter challenges in some implementations or specific application scenarios, failing to grasp the entire image’s global contextual details fully.
II-C2 Insufficiency in feature priority precision
Lacking a sophisticated method to discern the significance of diverse features, attention mechanisms might uniformly weigh all channel information. This approach risks overlooking vital features essential for precise classification, compromising the model’s effectiveness in identifying and leveraging distinctive patterns in the data.
II-C3 Inefficiency of Activation Functions
The reliance on ReLU within attention mechanisms can lead to the ”dying ReLU” issue, wherein neurons fail to activate over specific data ranges, stalling the learning progression. Such constraints significantly diminish the model’s adaptability during the training phase, adversely affecting the dynamics and overall efficacy of the attention-enhanced learning process.
II-C4 Lack of feature adaptability
Attention mechanisms may exhibit rigidity in adjusting and enhancing features non-linearly according to the unique characteristics of each input. This inflexibility can constrain the model’s ability to adapt and optimally leverage the diverse information within the data, thereby reducing the overall efficiency in feature utilization.
This study designed an innovative lightweight attention module to overcome the identified difficulties. This design incorporates adaptive average pooling, a two-layer fully connected network for channel recalibration, dual activation functions of LeakyReLU [10] and ReLU [11], and element-wise multiplication. Adaptive average pooling tackles the limitations in global feature extraction, whereas the two-layer fully connected network for channel recalibration optimizes feature priority precision. The combination of LeakyReLU and ReLU effectively addresses the inefficiencies of a single activation function, enhancing activation function efficiency, and element-wise multiplication improves feature adaptability. These pioneering approaches synergistically aim to significantly elevate model performance while mitigating computational resource requirements, providing a more efficient and adaptable solution for integrating attention mechanisms into breast cancer image classification models. Moreover, this paper presents a comparative analysis between the proposed DALAResNet50 model and the standard ResNet model, as illustrated in Table 2.
II-D Grad-CAM-based Visualization Techniques in Histopathology Image Analysis
Grad-CAM (Gradient-weighted Class Activation Mapping) is a widely used technique for visualizing the regions of an image that are most relevant to a neural network’s predictions. This technique has been extensively explored in the field of histopathology image analysis to improve model interpretability and provide insights into the decision-making process of deep learning models [59].
In histopathology, accurate interpretation of tissue images is crucial for diagnosing diseases such as cancer. Traditional Grad-CAM methods generate heatmaps by computing the gradients of the target class with respect to the feature maps of the last convolutional layer, highlighting the regions of the image that contribute most to the classification decision. This allows pathologists to understand which areas of the tissue image the model focuses on, aiding in verifying the model’s predictions.
However, conventional Grad-CAM has limitations, particularly in dealing with noisy activations and less distinct regions, which can hinder the clarity and precision of the visualizations. Various enhancements have been proposed to address these challenges. For instance, Guided Grad-CAM combines Grad-CAM with guided backpropagation to produce high-resolution class-discriminative visualizations, offering finer details of the important regions [60]. Additionally, Score-CAM removes the dependency on gradients by using the target class score, improving the heatmaps’ stability and robustness[61]. “
To further enhance the interpretability and effectiveness of visual explanations in histopathology image analysis, we chose to implement the Dynamic Threshold Grad-CAM (DT Grad-CAM) method. DT Grad-CAM addresses the limitations of conventional Grad-CAM by applying adaptive thresholding using Otsu’s method, which dynamically determines the optimal threshold for highlighting significant regions. This approach reduces noise and enhances the clarity of the visualizations, making it easier to identify and interpret the critical features that contribute to the model’s predictions. By focusing on the most relevant areas and minimizing the influence of less important regions, DT Grad-CAM provides more precise and interpretable heatmaps, which are particularly beneficial in clinical settings where accurate and transparent decision-making is essential.
III Methods
III-A Transfer ResNet50
The ResNet50 [9] model has a relatively complex structure, including a 50-layer deep network structure. The specific structure is shown in Table I. The primary module is made up of numerous residual blocks. Each block is equipped with multiple convolutional layers and distinctive identity maps. Uniquely, the ResNet50 model introduces the concept of residual learning. This innovation solves the problem of vanishing gradients in deep networks, thereby opening up the possibility of building deeper neural networks. This design breakthrough allowed for achieving promising performance. Therefore, the transfer learning strategy based on the pre-trained ResNet50 model is adopted in this paper, considering the efficient utilization of data and computing resources for model building. The benefit of transfer learning is the high-level feature representations learned by ResNet50 on large-scale datasets in medical image classification tasks. This allows for swiftly attaining superior classification outcomes, even with constrained data and computing capacities. As shown in Fig. 1, the proposed approach includes a histopathology breast cancer database and components for preprocessing and data augmentation, transfer learning based on ResNet50, and classification based on an improvised dual-activated fully connected layer. The proposed approach involves the following stages for breast cancer classification using histopathology images:
TABLE I: The ResNet50 Model
Module Name Output Size 50 Layers
Conv1 $112\times 112$ $7\times 7$, 64 stride=2
Conv2_x $56\times 56$ $\begin{Bmatrix}1\times 1,&64\\ 3\times 3&64\\ 1\times 1&256\end{Bmatrix}$ x3
Conv3_x $28\times 28$ $\begin{Bmatrix}1\times 1,&128\\ 3\times 3&128\\ 1\times 1&512\end{Bmatrix}$ x4
Conv4_x $14\times 14$ $\begin{Bmatrix}1\times 1,&256\\ 3\times 3&256\\ 1\times 1&1024\end{Bmatrix}$ x6
Conv5_x $7\times 7$ $\begin{Bmatrix}1\times 1,&512\\ 3\times 3&512\\ 1\times 1&2048\end{Bmatrix}$ x3
Pooling $1\times 1$ Average Pool, 1000-d FC.
Note: The term ”1000-d FC” in the pooling layer refers to a fully connected layer consisting of 1000 neurons.
Fig. 1: The framework of DALAResNet50 model
Fig. 2: Comparison of conventional ResNet50 and DALAResNet50 model: (a) ResNet50 model (b) DALAResNet50 model
Fig. 3: Visualization of DALAResNet50 model
40x
100x
200x
400x
Fig. 4: Images of breast cancer pathology at four magniations
III-B Dual-activated lightweight attention ResNet50
A new approach for automatic breast cancer classification is proposed in this research using a new version of the ResNet50 technique called Dual-activated lightweight attention ResNet50 (DALAResNet50). The DALAResNet50 model is methodically structured into three core stages: data preprocessing and augmentation, transfer learning based on ResNet50 architecture, and DALAResNet50 model optimization, as detailed below:
III-B1 Data preprocessing and augmentation
Initially, all breast cancer histopathology images are resized to 224x224 pixels, aligning with the input specifications for convolutional neural networks. A normalization process follows, adjusting pixel values to a range between [-1, 1] to expedite model training and augment the model’s generalization capability with novel data. Data augmentation strategies like random rotation and horizontal flipping bolster the model’s robustness. These strategies mimic potential real-world image alterations, improving the model’s resilience to subtle image variations.
III-B2 Transfer learning based on ResNet50
At this stage, capitalizes on a pre-trained ResNet50 framework, incorporating a custom-designed attention module within its fourth residual block. This module aims to pinpoint small-scale cell nuclei characteristic of breast cancer and glean pertinent semantic information. This strategic enhancement significantly boosts the model’s ability to learn and discern features.
III-B3 DALAResNet50 model optimization
During the optimization phase of the DALAResNet50 model, a sequential focus is placed on pooling layers, activation layers, and nonlinear feature adjustments to enhance the model’s performance and adaptability. The related steps are as follows:
  1. 1.
    Initially, to address the issue of spatial variations in input feature maps, traditional pooling layers are replaced with adaptive average pooling. This adjustment optimizes the network’s parameter structure, enabling the model to process global contextual information more effectively and accurately. The introduction of adaptive average pooling not only simplifies the network’s complexity but also enhances the model’s ability to capture important information from different parts of the image, laying a solid foundation for subsequent feature processing. Specifically, the adaptive average pooling operation is defined as follows:
    $y_{i}=\frac{1}{H_{i}\times W_{i}}\sum_{h=1}^{H_{i}}\sum_{w=1}^{W_{i}}x_{i,h,w}$ (1)
    where $y_{i}$ is the output of the adaptive average pooling layer for the $i$-th feature map, $H_{i}$ and $W_{i}$ are the height and width of the $i$-th feature map, respectively, and $x_{i,h,w}$ represents the value at position $(h,w)$ in the $i$-th feature map. This approach establishes a solid foundation for subsequent feature processing, allowing the model to effectively capture and utilize global contextual information from the input images.
  2. 2.
    Next, the model employs a dual activation function strategy in the activation layer, starting with the LeakyReLU activation function [10] in the initial phase. This function is chosen for its unique capability to maintain a slight yet nonzero gradient when the unit is inactive, effectively combating the problem of vanishing gradients and significantly boosting the training efficiency of deep neural networks. The LeakyReLU activation function is defined as follows:
    $f(x)=\begin{cases}x&\text{if }x>0\\ \alpha x&\text{otherwise}\end{cases}$ (2)
    where $f(x)$ represents the LeakyReLU function and $\alpha$ is a small constant (typically 0.01) that allows a small gradient when the input $x$ is negative. Following this, the ReLU activation function [11] is applied in the subsequent fully connected layer to optimize model performance further and enhance the representation of image features. The ReLU activation function is defined as:
    $f(x)=\max(0,x)$ (3)
    where (f (x)) is the ReLU function, which outputs (x) if (x) is positive and 0 otherwise. The careful combination of LeakyReLU and ReLU activation functions dramatically improves the model’s ability to process and learn complex image characteristics, providing strong support for accurately classifying breast cancer images. By employing LeakyReLU in the initial phase, the model effectively addresses the vanishing gradient problem, while the subsequent use of ReLU enhances the network’s nonlinearity and expressive power, thereby improving its overall performance and robustness.
  3. 3.
    Finally, the introduction of element-wise multiplication operations further enhances the model’s feature adaptability, allowing for dynamic, nonlinear adjustments of features based on specific inputs. This step significantly increases the network’s flexibility and information utilization, enabling the model to adapt its processing flow according to the particular circumstances of the input features. This approach notably improves the model’s feature processing capability, leading to higher accuracy and efficiency in classifying breast cancer images. The element-wise multiplication operation is defined as follows:
    $z_{i}=x_{i}\cdot y_{i}$ (4)
    where $z_{i}$ is the result of the element-wise multiplication of the $i$-th feature map, and $x_{i}$ and $y_{i}$ are the corresponding elements of the feature maps being multiplied. This operation allows the model to adjust the features dynamically, enhancing its ability to capture complex patterns and interactions within the data.
Fig. 2 illustrates the flowcharts of the standard ResNet50 approach and the DALAResNet50 approach, respectively. These flowcharts highlight the newly added features in the DALAResNet50. Fig. 3 depicts the visualization of the DALAResNet50 approach.
IV Experiment
IV-A Dataset
IV-A1 BreakHis Dataset
The BreakHis dataset, utilized in this study focusing on breast cancer histopathology images, was generously provided by scholar Spanha [38]. This dataset comprises histopathology samples of breast cancer tumors obtained from 82 patients. The samples were sectioned under a microscope and then imaged at different magnifications (40×, 100×, 200×, 400×) to capture intricate details. Experienced medical professionals meticulously interpreted and identified the images. The dataset consists of a total of 7,909 breast cancer pathological images, comprising 2,429 benign tumor images and 5,429 malignant tumor images. Each image has dimensions of 700 × 460 pixels and is stored in PNG format with three color channels. The detailed dataset statistics in Table 2 cover sample size, class distribution, etc. This multi-angle data resource provides a solid foundation for effective image classification and analysis research. In addition, to more intuitively display the breast cancer tissue details under different magnifications, Fig. 4 presents these pathological images as a practical example.
Since the BreakHis dataset is publicly available, this study does not require ethical approval. Additionally, publicly available datasets are typically subject to stringent quality control measures; therefore, this study does not consider image quality issues separately.
IV-A2 BACH Dataset
The BACH (Breast Cancer Histology Images) dataset [63] is a significant breast cancer histopathology image dataset provided by the ICIAR2018 Grand Challenge. This dataset includes 400 HE-stained histopathology images of breast cancer tissue, divided into four categories: Normal, Benign, In-situ carcinoma, and Invasive carcinoma, with 100 images in each category. All images have a resolution of 2048 × 1536 pixels, As shown in Table 3. The BACH dataset is widely used for training and testing machine learning and deep learning models to develop and evaluate breast cancer detection and classification algorithms. It provides high-quality annotation information, making it an essential resource in breast cancer research.
IV-A3 Mini-DDSM Dataset
The Mini-DDSM (Mini Digital Database for Screening Mammography) [62] is a reduced version of the DDSM dataset, primarily used for research and development in breast cancer screening. This dataset contains approximately 2000 mammography images, as shown in Table 4. a subset of the DDSM database that retains critical pathological features. The photos in the Mini-DDSM dataset are annotated with detailed information about lesions, including the location of masses and calcifications and their pathological diagnoses.
TABLE II: BreakHis dataset: benign and malignant image distribution across magnification factors
Class Magnification Factor
40X 100X 200X 400X
Benign 625 644 623 588
Malignant 1370 1437 1390 1232
Total No of Images 1995 2081 2013 1820
TABLE III: BACH Dataset Details
Category Number of Images Image Size
Normal 100 2048 × 1536
Benign 100 2048 × 1536
In-situ carcinoma 100 2048 × 1536
Invasive carcinoma 100 2048 × 1536
Total 400 2048 × 1536
TABLE IV: Detailed Information of the Mini-DDSM Dataset
Category Number of Images Image Size / Format
Normal Images 500 Variable
Benign Images 700 Variable
Malignant Images 800 Variable
Total Images 2000 Variable
IV-B Experimental environment
The experiments were carried out on an NVIDIA RTX A4000 GPU, utilizing the PyTorch 2.0.1 framework paired with Python 3.10.
IV-C Experiment implementation
This research utilizes a ratio of 6:1:3 to divide the processed dataset into training, verification, and test sets, as presented in Table 4. The training set is used for model training, during which parameter learning occurs. The verification set is employed to assess the model’s generalization ability. Continuous verification and parameter fine-tuning ensure superior model performance, with the optimal model pre-trained during verification. The test set measures the model’s recognition rate and generalization performance. Before model training, the training data is shuffled to enhance the model’s robustness and generalization capabilities.
In deep learning, the configuration of various parameters is paramount. Notably, the learning rate plays a crucial role; an excessively high learning rate may result in the loss value exploding, while a meager learning rate may lead to overfitting and reduce the model’s convergence speed. In previous studies, the initial learning rate is usually set in the range of 0.1 to 0.0001. Considering the relatively small size of the BreakHis medical imaging dataset (containing 7909 images) and the uneven distribution of the data, a lower learning rate of 0.0001 was chosen for this study so that the model could adjust the weights more accurately. This is important for dealing with small and imbalanced datasets to ensure that all categories are given proper attention. In addition, setting the training batch size and the number of iterations is significant. In addition, setting the training batch size and the number of iterations is substantial. Previous studies usually set the batch size between 32 and 128. Considering the GPU’s 16GB video memory limit, the lowest batch size of 32 was chosen for this study. As for the number of iterations, it is also usually set between 50 and 128. Given the size of the BreakHis dataset, this study limited the number of iterations to 50, aiming to balance training effectiveness and computational efficiency. Finally, this study also specifically considered the dropout rate setting. Previous studies generally set the dropout rate between 0.25 and 0.5. In processing the imbalanced BreakHis dataset, a lower dropout rate helps ensure that the categories with smaller numbers are sufficiently learned, thus improving the model’s recognition accuracy and generalization ability. Therefore, this study chose a lower dropout rate of 0.25.
The method proposed in this study and the benchmark model illustrated in Table 6 apply the above settings to ensure the experiments’ fairness and the comparability of the results. These carefully selected parameters provide a solid foundation for the experimental design and ensure the reliability and efficiency of the model throughout the study; the parameter settings are shown in Table 5.
TABLE V: Performance of Network Models on BreakHis Dataset (Two-Class Classification)
Network Models 40X 100X
Accuracy F1 IBA GMean Accuracy F1 IBA GMean
SEResNet50 [39] 0.883 0.779/0.883 0.860 0.860 0.850 0.720/0.849 0.813 0.812
DensNet121 [40] 0.801 0.639/0.800 0.763 0.762 0.811 0.654/0.809 0.755 0.753
VGG16 [41] 0.902 0.813/0.902 0.877 0.877 0.850 0.724/0.851 0.830 0.829
VGG16Inception [42] 0.977 0.955/0.977 0.965 0.965 0.971 0.943/0.971 0.964 0.964
ViT [43] 0.835 0.693/0.833 0.769 0.767 0.834 0.694 /0.833 0.788 0.787
Swin-Transformer [44] 0.843 0.711/0.844 0.819 0.818 0.877 0.767/0.876 0.836 0.835
Dinov2_Vitb14 [45] 0.873 0.762/0.873 0.843 0.843 0.869 0.753/0.868 0.835 0.834
RseNet50 [9] 0.927 0.860/0.927 0.914 0.914 0.944 0.891/0.944 0.922 0.922
DALAResNet50 0.985 0.970/0.985 0.980 0.980 0.987 0.974/0.987 0.985 0.985
Network Models 200X 400X
Accuracy F1 IBA GMean Accuracy F1 IBA GMean
SEResNet50 [39] 0.906 0.820/0.906 0.876 0.875 0.876 0.770/0.878 0.864 0.863
DensNet121 [40] 0.812 0.659/0.812 0.775 0.774 0.779 0.603/0.777 0.737 0.736
VGG16 [41] 0.851 0.722/0.850 0.812 0.811 0.848 0.718/0.848 0.817 0.817
VGG16Inception [42] 0.954 0.910/0.954 0.938 0.938 0.889 0.788/0.888 0.857 0.856
ViT [43] 0.841 0.705/0.840 0.801 0.800 0.832 0.691/0.831 0.803 0.803
Swin-Transformer [44] 0.876 0.766/0.875 0.838 0.837 0.846 0.724/0.851 0.791 0.789
Dinov2_Vitb14 [45] 0.853 0.726/0.852 0.802 0.800 0.885 0.782/0.885 0.840 0.839
RseNet50 [9] 0.960 0.922/0.960 0.947 0.947 0.925 0.856/0.925 0.918 0.918
DALAResNet50 0.979 0.958/0.979 0.98 0.98 0.943 0.891/0.944 0.945 0.945
Note: In the table, the DenseNet-121 model indicates that employed a pre-trained DenseNet-121 model for transfer learning.
TABLE VI: Performance of Network Models on BACH Dataset (Four-Class Classification)
Network Models Accuracy F1 IBA GMean
SEResNet50 [39] 0.770 0.743/0.77 0.583 0.553
DensNet121 [40] 0.703, 0.680/0.703 0.458 0.438
VGG16 [41] 0.788 0.781 /0.788 0.652 0.637
VGG16Inception [42] 0.727 0.864/0.727 0.650 0.654
ViT [43] 0.746 0.734/ 0.745 0.580 0.555
Swin-Transformer [44] 0.748 0.556/0.745 0.500 0.436
Dinov2_Vitb14 [45] 0.856 0.730/0.855 0.500 0.353
ResNet50 [9] 0.855 0.838 /0.855 0.664 0.635
DALAResNet50 0.891 0.891/0.885 0.748 0.734
TABLE VII: Performance of Network Models on Mini-DDSM Dataset (Three-Class Classification)
Network Models Accuracy F1 IBA GMean
SEResNet50 [39] 0.7006 0.699/0.701 0.709 0.768
DensNet121 [40] 0.6301 , 0.627/0.63 0.642 0.716
VGG16 [41] 0.6607 0.66 /0.661 0.672 0.736
VGG16Inception [42] 0.7247 0.728/0.725 0.728 0.786
ViT [43] 0.7515 0.72/ 0.752 0.580 0.555
Swin-Transformer [44] 0.5110 0.51/ 0.511 0.52 0.619
Dinov2_Vitb14 [45] 0.667 0.665/0.667 0.678 0.742
ResNet50 [9] 0.6686 0.672 /0.669 0.677 0.74
DALAResNet50 0.8070 0.806/0.807 0.818 0.851
TABLE VIII: Parameters Count and Convergence Time of Network Models on BreakHis Dataset
Network Models Convergence Time (seconds) Total Parameters Average Time
40X 100X 200X 400X
SEResNet50 [39] 2696.22 2814.30 2711.89 2356.18 27073090 2644.65
DenseNet121 [40] 2352.46 2643.79 2423.36 2078.34 9028970 2347.49
VGG16 [41] 3212.95 3182.54 3011.53 2640.19 134268738 3011.80
VGG16Inception [42] 3074.99 3253.50 3303.12 2682.61 38139938 3078.56
ViT [43] 2932.11 3022.54 3013.86 2691.16 51375106 2914.92
Swin-Transformer [44] 3020.11 3132.42 2940.98 2592.96 86745274 2921.62
Dinov2_Vitb14 [45] 2895.02 2303.25 2245.98 2017.15 86582018 2365.35
ResNet50 [9] 2920.53 2933.26 2307.94 2097.33 21797672 2564.77
DALAResNet50 2133.02 2319.50 1990.58 1896.82 25611842 2084.98
Note: In the table, the DenseNet-121 model indicates that a pre-trained DenseNet-121 model was employed for transfer learning.
TABLE IX: Parameters Count and Convergence Time of Network Models on BACH Dataset
Network Models Total Parameters Average Time
SEResNet50 [39] 27074116 2921.67
DensNet121 [40] 9029996 3275.27
VGG16 [41] 134276934 3303.98
VGG16Inception [42] 38148134 3432.98
ViT [43] 51385352 2872.36
Swin-Transformer [44] 86753470 3488.72
Dinov2_Vitb14 [45] 86583556 2966.12
ResNet50 [9] 21805868 3723.49
DALAResNet50 24042692 2475.49
TABLE X: Parameters Count and Convergence Time of Network Models on Mini-DDSM Dataset
Network Models Total Parameters Average Time
SEResNet50 [39] 27073603 7272.75
DensNet121 [40] 9029483 6384.11
VGG16 [41] 134272835 7589.96
VGG16Inception [42] 38144035 8089.23
ViT [43] 51376131 6555.35
Swin-Transformer [44] 86746299 9715.44
Dinov2_Vitb14 [45] 86582787 6432.41
ResNet50 [9] 21797672 8268.68
DALAResNet50 24040643 5990.31
IV-D Baseline selection
Table 6 presents a comparative analysis of various deep learning methods, encompassing widely adopted approaches and the novel strategy introduced in this paper. The following section briefly outlines several of the models included in the comparison:
SEResNet50 [39]: Hu et al. proposed this model in 2018. The key feature of this model is integrating the ”squeeze and excitation” attention mechanism, which aims to improve the network’s capabilities.
DenseNet-121 [40]: Huang et al. proposed this model in 2017 by introducing dense connections (dense blocks) to transfer and reuse information between network layers, thus improving the efficiency of feature transmission and reproduction. This study employs transfer learning with DenseNet-121, which is utilized for classifying breast cancer. The advantage of using a pre-trained DenseNet is that it can adapt to small sample data.
VGG16 [41]: This model is from Simonyan and Zisserman and was released in 2014. It is known for its clean structure and thoughtful design philosophy and is widely used for image classification and feature extraction tasks, relying mainly on deeply stacked convolutional layers.
VGG16Inception [42]: The model is presented by M. Saini and S. Susan. The article delves into a deep learning model called ”VGGIN-Net,” which is mainly applied to analyze imbalanced breast cancer datasets. The study introduces a new deep learning framework that incorporates transfer learning to address the problem of category imbalance in biomedical datasets, applied explicitly to the BreakHis breast cancer dataset. The model uses a pre-trained network to migrate learned weights from one domain to another to reduce training time and computational cost.
ViT (Vision Transformer) [43]: Proposed by Dosovitskiy et al. in 2020, it has gained widespread attention by successfully applying the Transformer architecture to image recognition tasks. This model has shown great strength in image recognition.
Swin-Transformer [44]: Liu et al. created this model in 2021, and its performance in various vision tasks is outstanding. This model is unique in that it solves the computational and memory problems that can be encountered when processing large images.
Dinov2_Vitb14 [45]: Proposed in 2023 by a team of researchers at Facebook, these models contain a pre-training dataset of 142 million unlabelled images to obtain efficient visual properties demonstrating general solid performance.
Each of the above models has uniqueness and excellent image classification and vision tasks. To emphasize the effectiveness of the proposed approach, an in-depth comparison of all these models has been conducted, achieving significant improvements in various evaluation metrics.
IV-E Experiment metrics
Used four metrics, Accuracy [46, 47, 48], F1 Score [49], Imbalance Balanced Accuracy (IBA) [50], and Geometric Mean (GMean) [51], to assess the performance of the model holistically. Below is an introduction to these metrics:
  1. 1.
    Accuracy is one of the most intuitive and commonly used metrics for evaluating the overall performance of a model. It calculates the proportion of all samples (whether positive or negative) that were predicted correctly relative to the total number of samples [46, 47, 48].
  2. 2.
    The F1 Score is a crucial metric that combines precision and recall, providing a comprehensive measure of a model’s ability to classify positive and negative instances. This is especially useful for imbalanced datasets where positive and negative samples differ significantly [52].
  3. 3.
    IBA is a metric for evaluating the performance of classification models on imbalanced datasets, adjusting standard accuracy to more fairly account for the model’s ability to identify both majority and minority classes, ensuring the performance assessment reflects classification effectiveness across all categories. [53].
  4. 4.
    GMean is a metric that calculates the geometric mean of individual class accuracies, indicating a model’s capability to address category imbalance. It effectively equalizes the performance assessment for both dominant and infrequent categories, emphasizing the model’s comprehensive efficiency in handling variances in class distribution. This metric is in scenarios where class imbalance could significantly skew the performance metrics, ensuring that the model’s ability to classify less represented classes correctly is adequately considered in its overall [54].
Tables 5, 6, and 7 present the model’s performance on the binary classification BrekHis dataset, the four-class classification BACH dataset, and the three-class classification Mini-DDSM dataset. In these results, the model consistently outperforms nine other popular networks in terms of accuracy, F1 score, IBA, and GMean. This consistency highlights the significant advantages of the network and underscores its potential in medical image classification tasks.
The model was compared with eight other models in terms of parameter count and convergence speed on the binary classification BrekHe dataset, the four-class classification BACH dataset, and the three-class classification Mini-DDSM dataset, as shown in Tables 8, 9, and 10. Although DenseNet-121 has the fewest parameters, the proposed model achieved the shortest convergence time. This further reinforces the efficiency and robust potential of the proposed model in medical image classification, highlighting its superior performance and effectiveness.
IV-F Ablation study
In this ablation study, we compared the performance of the DALAResNet50 model with the standard ResNet50 model on the binary classification BrekHis dataset at 40x, 100x, 200x, and 400x resolutions, as shown in Table 5 above. We also provided the confusion matrices for both models on the test set, as illustrated in Fig. 5. The results indicate that the DALAResNet50 model demonstrates higher sensitivity and specificity at all magnification levels, significantly reducing the number of false positives and false negatives. These findings suggest that the DALAResNet50 model outperforms the standard ResNet50 model in terms of classification accuracy and reliability.
V DT Grad-CAM
The interpretability of deep learning models is crucial, particularly in areas such as breast cancer diagnosis. Grad-CAM is a popular technique for visualizing class-specific regions of interest in images. This article introduces an enhancement called DT Grad-CAM (Dynamic Threshold Grad-CAM), which improves Grad-CAM by applying adaptive thresholds using the Otsu method, enhancing the visualizations’ clarity and relevance.
V-A DT Grad-CAM method
The DT Grad-CAM method enhances the traditional Grad-CAM approach through several steps. Each step improves the visualization of class-specific regions, making the results more interpretable and relevant for diagnostic purposes.
1. Grad-CAM Calculation
Grad-CAM uses the following formula to compute the gradient-weighted feature map:
$L_{c}^{\text{Grad-CAM}}=\text{ReLU}\left(\sum_{k}\alpha_{k}^{c}A^{k}\right)$ (5)
where:
  • $L_{c}^{\text{Grad-CAM}}$ is the Grad-CAM heatmap for class $c$.
  • $A^{k}$ is the $k$-th feature map of the convolutional layer.
  • $\alpha_{k}^{c}$ is the weight for feature map $A^{k}$, defined as the global average pooling of the gradients of class $c$ with respect to $A^{k}$:
$\alpha_{k}^{c}=\frac{1}{Z}\sum_{i}\sum{j}\frac{\partial y^{c}}{\partial A_{ij}^{k}}$ (6)
where:
  • $y^{c}$ is the score for class $c$.
  • $\frac{\partial y^{c}}{\partial A_{ij}^{k}}$ is the gradient of the score for class $c$ with respect to the feature map $A^{k}$ at location $(i,j)$.
    40x
    100x
    200x
    400x
    Figure 5: Comparison of the confusion matrices of the DALAResNet50 model and the standard ResNet50 model.
    40X- Original
    40X-DTGrad-CAM
    40X- Grad-CAM
    100X- Original
    100X-DTGrad-CAM
    100X- Grad-CAM
    200X- Original
    200X-DTGrad-CAM
    200X- Grad-CAM
    400X- Original
    400X-DTGrad-CAM
    400X- Grad-CAM
    Figure 6: Comparison of Original Histopathology Images with DT Grad-CAM and Grad-CAM Visualizations at Different Magnifications
  • $Z$ is the total number of elements in the feature map (i.e., width times height).
This step leverages the gradient information to highlight the most important features contributing to the class prediction, providing an intuitive visual explanation of the model’s decision-making process.
2. Otsu Threshold Calculation
After generating the Grad-CAM heatmap, DT Grad-CAM applies Otsu’s method to determine the optimal threshold value. Otsu’s method calculates the threshold $T$ that maximizes the between-class variance:
$\sigma_{B}^{2}(T)=\omega_{1}(T)\omega_{2}(T)[\mu_{1}(T)-\mu_{2}(T)]^{2}$ (7)
where:
  • $\sigma_{B}^{2}(T)$ is the between-class variance for threshold $T$.
  • $\omega_{1}(T)$ and $\omega_{2}(T)$ are the probabilities of foreground and background pixels at threshold $T$, respectively.
  • $\mu_{1}(T)$ and $\mu_{2}(T)$ are the means of the foreground and background pixels at threshold $T$, respectively.
By iterating through all possible thresholds $T$, the one that maximizes $\sigma_{B}^{2}(T)$ is selected as the Otsu threshold.
Using Otsu’s method allows for adaptive thresholding based on the image’s specific characteristics, ensuring that the most relevant regions are highlighted dynamically rather than relying on a fixed threshold.
3. Thresholding
The Otsu threshold is then used to binarize the Grad-CAM heatmap:
$L_{c}^{\text{Thresh}}(x,y)=\begin{cases}L_{c}^{\text{Grad-CAM}}(x,y)&\text{if }L_{c}^{\text{Grad-CAM}}(x,y)>T\\ 0&\text{otherwise}\end{cases}$ (8)
where:
  • $L_{c}^{\text{Thresh}}$ is the thresholded Grad-CAM heatmap.
  • $T$ is the optimal threshold determined by Otsu’s method.
This thresholding step ensures that only the most significant regions in the Grad-CAM heatmap are highlighted, improving the visualization’s clarity and focus by removing less relevant areas.
V-B Comparison between DT Grad-CAM and Grad-CAM Visualization
Fig. 6 illustrates a side-by-side comparison of original images, DT Grad-CAM visualizations, and traditional Grad-CAM visualizations at various magnifications (40X, 100X, 200X, and 400X). Each row shows the original image, the DT Grad-CAM visualization, and the Grad-CAM visualization, respectively. The DT Grad-CAM method consistently provides clearer and more focused representations of significant regions across all magnifications. By applying adaptive thresholding using Otsu’s method, DT Grad-CAM reduces noise and enhances the visibility of critical areas, which is particularly evident when comparing the middle columns (DT Grad-CAM) with the right columns (Grad-CAM). This adaptive approach results in sharper and more interpretable visualizations, highlighting the relevant features more distinctly and minimizing the influence of irrelevant regions. Consequently, DT Grad-CAM offers superior clarity and relevance, making it a valuable tool for improving model interpretability, especially in diagnostic applications such as breast cancer imaging.
Table 11 provides a detailed comparison between DT Grad-CAM and Grad-CAM, highlighting the key differences and advantages of DT Grad-CAM. The adaptive thresholding in DT Grad-CAM enhances the clarity and relevance of the visualizations by minimizing noise and focusing on the most significant features. The clarity of visualization is significantly improved in DT Grad-CAM by isolating the key areas of interest and reducing the influence of less relevant regions. This enhancement in clarity directly translates to better interpretability, as DT Grad-CAM offers a sharper and more meaningful highlight of critical regions, making it easier for users to understand and interpret the model’s focus areas. Noise reduction is another major advantage of DT Grad-CAM, as it effectively reduces noise in the visualization, ensuring that only the most pertinent areas are highlighted. This is particularly useful in diagnostic applications, such as breast cancer imaging, where precise and clear visualization of important regions is crucial for accurate analysis. In contrast, traditional Grad-CAM lacks these adaptive enhancements, often resulting in more noise and less precise highlighting of important regions, which can obscure the interpretation and impact the accuracy of diagnostic applications.
TABLE XI: Comparison between DT Grad-CAM and Grad-CAM
Feature DT Grad-CAM Grad-CAM
Adaptive Thresholding Utilizes Otsu’s method to dynamically determine the optimal threshold, enhancing the clarity and relevance of the highlighted regions by minimizing noise and focusing on the most significant features. Does not employ adaptive thresholding, often leading to more noise and less precise highlighting of important regions.
Clarity of Visualization Provides clearer and more focused visualizations by isolating the key areas of interest and reducing the influence of less relevant regions. Visualizations may be more dispersed and less distinct, making it harder to interpret the significant regions accurately.
Interpretability Enhances model interpretability by offering a sharper and more meaningful highlight of critical regions, making it easier for users to understand and interpret the model’s focus areas. Provides a basic level of interpretability but can be less effective in highlighting the most relevant features clearly.
Noise Reduction Effectively reduces noise in the visualization, ensuring that only the most pertinent areas are highlighted, thus improving the overall quality of the heatmap. Higher levels of noise can obscure the interpretation, as less important regions may also be highlighted.
Relevance in Diagnostic Applications Particularly useful in diagnostic applications, such as breast cancer imaging, where precise and clear visualization of important regions is crucial for accurate analysis. Useful but may lack the precision required for detailed diagnostic purposes, potentially impacting the accuracy of interpretations.
VI Performance evaluation and visualization
To effectively monitor the model’s performance and gain insights into the training process, it is essential to concentrate on various crucial metrics, such as training loss, validation loss, training accuracy, and validation accuracy. Visualizing these metrics offers a comprehensive view of the model’s training behaviour. An in-depth analysis of these key metrics and their implications is outlined below:
VI-A Loss analysis
This study deployed a DALAResNet50 model for the classification of breast cancer. Critical metrics for assessing model performance and understanding the training dynamics include training loss and validation loss [55, 56, 57]. These metrics are crucial for evaluating the model’s effectiveness, and their graphical representation offers a detailed insight into the training progression.
Fig. 7 illustrates the fluctuations in training and validation losses across varying numbers of training iterations at different image magnifications (40X, 100X, 200X, and 400X). The x-axis represents the number of training iterations, while the y-axis reflects the loss values. Line plots demonstrate the trends in training and validation losses, with error bars on each point indicating the standard deviation of these values.
Training and validation losses consistently decline, indicating the model’s successful learning and adaptation to the dataset. This downward trend is encouraging, as it signals ongoing improvement in task performance. In the later training phases, the validation loss stabilizes, suggesting that the model avoids overfitting on the validation dataset and exhibits robust generalization capabilities to unseen data.
VI-B Accuracy analysis
In Fig.8, the progression of training and validation accuracy across various image resolutions (40X, 100X, 200X, and 400X) is illustrated [58]. The horizontal axis indicates the number of training iterations, whereas the vertical axis displays accuracy levels. A line graph is utilized to depict the trend in accuracy, with error bars included to illustrate the standard deviation of the accuracy figures.
After meticulous analysis, the following outcomes have been derived: Training and validation accuracy undergo a gradual ascent, showcasing the model’s persistent learning and adaptation throughout the training process. This upward trajectory suggests that the model is advancing in comprehending and classifying the data. Additionally, the validation accuracy maintains stability, especially in the later stages of training, implying that the model avoids overfitting to the training data and can generalize effectively to unseen data. These findings are pivotal in comprehending the model’s performance and can facilitate further optimization for enhanced breast cancer classification results.
VII DALAResNet50’s Future Therapeutic Impact and Integration into Existing Diagnostic Processes
DALAResNet50 model promises to significantly impact the future of breast cancer treatment by enhancing diagnostic accuracy and providing clear, interpretable visualizations through the integration of DT Grad-CAM. As healthcare systems increasingly adopt artificial intelligence (AI) technologies, This model can streamline diagnostic workflows, reduce diagnostic errors, and facilitate earlier detection of breast cancer, which is crucial for improving patient outcomes.
Incorporating DALAResNet50 into existing diagnostic processes can enhance the efficiency and accuracy of histopathology evaluations. Pathologists can leverage the model’s high classification accuracy to identify cancerous tissues more confidently, thus supporting more accurate and timely diagnoses. The use of DT Grad-CAM further aids pathologists by providing visual explanations for the model’s predictions, highlighting critical regions within histopathology images, and ensuring transparency in the diagnostic process.
Moreover, the DALAResNet50 model can serve as a valuable second-opinion tool, assisting pathologists in confirming their initial diagnoses and potentially identifying cases that may require further review. This integration can lead to a more robust diagnostic process, where AI augments human expertise, reducing the cognitive load on pathologists and minimizing the risk of oversight.
As the model continues to be refined and validated with larger and more diverse datasets, its application can extend beyond breast cancer to other types of histopathological analyses. This scalability and adaptability make DALAResNet50 a versatile tool in the broader field of medical diagnostics.
In the future, integrating DALAResNet50 with electronic health record (EHR) systems can further streamline patient management by providing seamless access to diagnostic results and AI-generated insights. This integration would facilitate multidisciplinary collaboration among healthcare providers, ensuring that treatment plans are informed by the most accurate and comprehensive diagnostic information available.
Overall, the DALAResNet50 model, combined with DT Grad-CAM’s interpretability, holds great promise for transforming breast
40X
100X
200X
400X
Fig. 7: Training and validation accuracy with error bars
40X
100X
200X
400X
Fig. 8: Training and validation loss with error bars
cancer diagnosis and treatment. Its integration into existing diagnostic workflows can lead to more precise, reliable, and efficient medical care, ultimately enhancing patient outcomes and advancing the field of histopathology.
VIII Summary and conclusions
This paper proposed a DALAResNet50 model for classifying medical histopathology tissue images, mainly when dealing with limited-scale imbalanced breast cancer histopathology image datasets. The ResNet50 model incorporates the lightweight attention mechanism to optimize the pre-trained deep network layer. The performance evaluation includes metrics such as accuracy, F1 score, IBA, and GMean, which allow for a comprehensive assessment of this paper’s model performance compared to existing techniques. The experimental results highlight significant advantages, particularly when addressing the challenges of imbalanced breast cancer datasets. The model demonstrates improved robustness, generalization, faster convergence during training, and effective prevention of overfitting. These findings provide a solid basis for enhancing classification accuracy. Furthermore, this study introduces the DT Grad-CAM method to address the interpretability challenges often associated with deep learning models. DT Grad-CAM enhances the traditional Grad-CAM approach by applying adaptive thresholding using Otsu’s method, which provides clearer and more focused visualizations of significant regions. This method improves the clarity and relevance of the highlighted areas, making it easier for medical experts to interpret the model’s predictions and understand the critical features contributing to the classification. The integration of DT Grad-CAM with DALAResNet50 ensures high classification performance and enhanced interpretability, crucial for reliable and transparent breast cancer diagnosis. In future research, we plan to apply the model presented in this paper to a broader range of medical fields.